Weiss Manufacturing intends to increase capacity by overcoming a bottleneck operation by adding new equipment. Two vendors have presented proposals. The fixed costs are

$ 50 comma 000$50,000

for proposal A and

$ 70 comma 000$70,000

for proposal B. In addition to the proposed fixed costs from the two​ vendors, Weiss's management anticipates that they will have to spend

$ 12 comma 000$12,000

for installations to be completed. The variable cost is

$ 14.00$14.00

for A and

$ 12.00$12.00

for B. The revenue generated by each unit is

$ 20.00$20.00.

​a) The​ break-even point in dollars for the proposal by Vendor A​ =

​$nothing

​(round your response to the nearest whole​ number).
Answer #1

Vendor A

Let the point of indifference be X,

Total Expense = Total Revenue

Fixed Cost + Variable Cost = Selling price * Volume

50000 + 12000 + 14*X = 20*X

6*X = 62000

X = 10333.33 Units

Break even point in dollars = 20*10333.33

= $206667

Please provide up votes, if you are satisfied with the response. Comment in case of any issue with the same.